Opened 10 years ago

Closed 9 years ago

Last modified 3 years ago

#285 closed flight processing (fixed)

ET07/01, flight day 300b/2009, Afar south

Reported by: benj Owned by: knpa
Priority: alpha 5 Milestone: 2009 Data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by mark1)

Data location: ~arsf/arsf_data/2009/flight_data/ethiopia/ET07_01-2009_300b_Afar_south

Data arrived from ARSF via SATA disk Y on 4th Nov. 2009.

Scientific objective: /users/rsg/arsf/arsf_data/2008/ARSF_Applications/Ethiopia/ET07_01\ -\ Dr\ Wright.pdf

Priority: alpha-5

PI: Tim Wright

2009 Ethiopia campaign, reflown from 2007

Sensors:

  • Eagle (12/02/2010)
  • Hawk (12/02/2010)
  • Leica LIDAR (12/02/2010)
  • Digital Photography (26/02/2010)
  • TABI (to be processed by University of Athens)

Change History (22)

comment:1 Changed 10 years ago by benj

  • Description modified (diff)

Raw GPS base station data is in ~arsf/arsf_data/2009/flight_data/ethiopia/basestations.

comment:2 Changed 10 years ago by benj

  • Description modified (diff)

comment:3 Changed 10 years ago by mggr

  • Description modified (diff)

Raw TABI data acquired, put in archive and offered to Athens for download and processing.

comment:4 Changed 10 years ago by emca

  • Owner set to emca
  • Status changed from new to assigned

comment:5 Changed 10 years ago by emca

Eagle and Hawk SCT offsets per line

Line Eagle Hawk
1-0.08 -0.06
2-0.05 -0.04
3-0.04 -0.06
4+0.02 -0.07
5-0.01 -0.05
60.00 -0.05
7-0.03 -0.07
8-0.04 -0.04
90.00 -0.08
10-0.05 -0.07
11-0.02 -0.03
12+0.02 -0.01
13-0.05 -0.01
14+0.02 -0.07
15-0.06 -0.08

comment:6 Changed 10 years ago by emca

Amended logsheet to contain correct Project Code (was entered as ET07-04)

comment:7 Changed 10 years ago by emca

Processing complete.

Hyperspectral data ready for delivery check at:
/users/rsg/airborne/workspace/ET07_01-2009_300b_Afar_south/delivery/20100125/

LIDAR ready for delivery check at:
/users/rsg/airborne/workspace/ET07_01-2009_300b_Afar_south/delivery/20100126/

comment:8 Changed 10 years ago by mark1

Photograph delivery created and checked. Ready to deliver.

comment:9 Changed 10 years ago by mark1

LIDAR delivery check: minor corrections made to read_me, all ascii files checked with check_ascii_lidar.sh and viewed in grass. Data looks good.

DEM header resolution and Read_me disagree, this needs to be checked as to which is correct. Also need to add some info into readme about SRTM data etc.

comment:10 Changed 10 years ago by emca

Re: LIDAR
Readme has been corrected.
Data now ready for delivery.

comment:11 Changed 10 years ago by knpa

Hyperspectral has been delivery checked.

All OK apart from:

  • Contents list in readme doesn't include misc or bin folders and contents
  • Quality of Data section - says we processed our lev3s with SRTM DEM; was actually SRTM and lidar(?)
  • Geocorrection of Data section - should mention that we have included an SRTM/lidar DEM with the delivery.
  • Need to add -es misc/sphsep15lx.grd to example azgcorr command
  • Need to mention in Data Quality that Eagle is reguarly dropping scan lines
  • Hawk 01 has a large number of dropped scan lines at the start of the line - all bands

comment:12 Changed 10 years ago by emca

Hyperspectral:
Read_me edited for above changes

Note:

  • Misc directory not included as not required for delivery
  • Data was quality checked using SRTM DEM
  • -es NO used for sample azgcorr command


Data now ready for delivery

comment:13 Changed 10 years ago by emca

LIDAR and Hyperspectral sent to PI 12/02/2010

comment:14 Changed 10 years ago by mark1

  • Description modified (diff)

Digital Photography dispatched to PI on 26/02/2010

comment:15 Changed 10 years ago by benj

Processed TABI data was received from Pol back in January but we weren't happy with the accuracy. This data is in the workspace, it has not been dispatched to the PI but is the only processed TABI data we currently have for this project.

comment:16 Changed 9 years ago by knpa

  • Component changed from Processing: general to Archiving
  • Owner changed from emca to knpa

Appears to be r-synced (apart from Tabi).

Am going to proceed with archiving, despite the unresolved tabi issue, due to the need for disk space. Will exclude Tabi from archived tarball but keep what data we have for it in the repository.

comment:17 Changed 9 years ago by knpa

Missing ipas intermediate data, i.e. extract and gps folders.

comment:18 Changed 9 years ago by knpa

Finished preparing for archiving and have created tarball.
NEODC have been informed.

The Tabi data has not been included in the tarball and has been relocated to:

~arsf/arsf_data/2009/flight_data/ethiopia/tabi_ET07_01-2009_300b_Afar_south/

comment:19 Changed 9 years ago by knpa

Downloaded and backed up by NEODC.

Moved to archive disk.

comment:20 Changed 9 years ago by knpa

  • Resolution set to fixed
  • Status changed from assigned to closed

Deleted workspace copy.

Closing ticket.

comment:21 Changed 3 years ago by lah

Tidying

There is no reprocessed hyperspectral delivery, but there are reprocessed files, so I've created a new delivery out of these and added a readme. This will need a delivery check before archiving. I have not deleted any of the hyperspectral processing files yet - will wait until after the delivery is checked.

There was a problem getting the under/overflows from the mask files, so this is not included in the readme, but might indicate a problem with the files.

comment:22 Changed 3 years ago by asm

Archiving

Archived and it is available from CEDA. Re-processed hyperspectral data at http://browse.ceda.ac.uk/browse/neodc/arsf/2009/ET07_01/ET07_01-2009_300b_Afar_south/hyperspectral-reprocessed

Plymouth 12th May 2017

Note: See TracTickets for help on using tickets.